# Yuhui Li1,3, Fangyun Wei2, Chao Zhang1, Hongyang Zhang3,4

### v1

@@ Abstract
The sequential nature of modern LLMs makes them expensive and slow, and speculative sampling has proven to be an effective solution to this problem. Methods like EAGLE perform autoregression at the feature level, reusing top-layer features from the target model to achieve better results than vanilla speculative sampling. A growing trend in the LLM community is scaling up training data to improve model intelligence without increasing inference costs. However, we observe that scaling up data provides limited improvements for EAGLE. We identify that this limitation arises from EAGLE’s feature prediction constraints. In this paper, we introduce EAGLE-3, which abandons feature prediction in favor of direct token prediction and replaces reliance on top-layer features with multi-layer feature fusion via a technique named training-time test. These improvements significantly enhance performance and enable the draft model to fully benefit from scaling up training data. Our experiments include both chat models and reasoning models, evaluated on five tasks. The results show that EAGLE-3 achieves a speedup ratio up to 6.5x, with about 1.4x improvement over EAGLE-2. The code is available at https://github.com/SafeAILab/EAGLE.
@@
Yuhui Li1,3, Fangyun Wei2, Chao Zhang1, Hongyang Zhang3,4 1Peking University, 2Microsoft Research, 3University of Waterloo, 4Vector Institute. yuhui.li@stu.pku.edu.cn, fawe@microsoft.com c.zhang@pku.edu.cn, hongyang.zhang@uwaterloo.ca
@@ 1 Introduction
  • Improved inference acceleration: EAGLE-3, trained with approximately 8x more data than EAGLE, achieves a 1.4x latency speedup over EAGLE-2 at batch size 1. For large-batch inference, EAGLE-3 also surpasses the implementation of vLLM Kwon et al. (2023) in throughput, even at batch sizes up to 56. We expect larger data size would lead to further improved speedup ratio.
@@ 4.2 Ablation Study
4.3 Throughput
Speculative sampling algorithms like EAGLE-3 reduce memory accesses and lower latency during memory-bound decoding by leveraging redundant computational power. As batch sizes increase, this redundancy decreases, reducing the effectiveness of speculative sampling. We conducted a study on the impact of EAGLE-3 on throughput for large batch sizes based on vLLM, a widely used production-grade framework, and the results are shown in Table 3. EAGLE shows the maximum throughput improvement at a batch size of 24, while EAGLE-3 shows this at 56. This part of the experiment did not use the tree structure, and the maximum chain length was set to 2. The framework used is also different from the other sections of this paper and from the typical settings in the speculative sampling domain, so comparisons with other experimental results are not meaningful.
Table 3: Throughput improvement under different batch sizes, with vLLM without speculative sampling as the baseline (1.00x).

### v2

@@ Abstract
The sequential nature of modern LLMs makes them expensive and slow, and speculative sampling has proven to be an effective solution to this problem. Methods like EAGLE perform autoregression at the feature level, reusing top-layer features from the target model to achieve better results than vanilla speculative sampling. A growing trend in the LLM community is scaling up training data to improve model intelligence without increasing inference costs. However, we observe that scaling up data provides limited improvements for EAGLE. We identify that this limitation arises from EAGLE’s feature prediction constraints. In this paper, we introduce EAGLE-3, which abandons feature prediction in favor of direct token prediction and replaces reliance on top-layer features with multi-layer feature fusion via a technique named training-time test. These improvements significantly enhance performance and enable the draft model to fully benefit from scaling up training data. Our experiments include both chat models and reasoning models, evaluated on five tasks. The results show that EAGLE-3 achieves a speedup ratio up to 6.5x, with about 1.4x improvement over EAGLE-2. In the SGLang framework, EAGLE-3 achieves a 1.38x throughput improvement at a batch size of 64. The code is available at https://github.com/SafeAILab/EAGLE.
@@
Yuhui Li1,3, Fangyun Wei2, Chao Zhang1, Hongyang Zhang3,4 1Peking University 2Microsoft Research 3University of Waterloo 4Vector Institute yuhui.li@stu.pku.edu.cn, fawe@microsoft.com c.zhang@pku.edu.cn, hongyang.zhang@uwaterloo.ca
@@ 1 Introduction
  • Improved inference acceleration: EAGLE-3, trained with approximately 8x more data than EAGLE, achieves a 1.4x latency speedup over EAGLE-2 at batch size 1. Speculative sampling is often thought to reduce throughput at large batch sizes. However, in SGLang Zheng et al. (2024), a production-grade framework, EAGLE-3 improves throughput by 40% at a batch size of 64. We expect larger data size would lead to further improved speedup ratio.
@@ 4.2 Ablation Study
4.3 EAGLE-3 in SGLang
Speculative sampling algorithms like EAGLE-3 reduce memory accesses and lower latency during memory-bound decoding by leveraging redundant computational power. As batch sizes increase, this redundancy decreases, reducing the effectiveness of speculative sampling. Efficiency improvements are more challenging in highly optimized production-grade frameworks. The performance of EAGLE-3 for large batches on a single H100 GPU and LLaMA-Instruct 3.1 8B in the SGLang v0.4.4 environment Zheng et al. (2024) was evaluated by the SGLang team, shown in Table 3. This part of the experiment did not use the tree structure, the chain length was set to 3, and the testing dataset was MT-Bench. EAGLE reduces throughput at a batch size of 24, whereas EAGLE-3 still achieves a 38% throughput improvement at a batch size of 64.
Table 3: Throughput improvement under different batch sizes on H100 and LLaMA-Instruct 3.1 8B for the MT-Bench dataset, with SGLang without speculative sampling as the baseline (1.00x). The experiments were conducted by the SGLang team.
Batch size 2 4 8 16 24 32 48 56 64
EAGLE 1.40x 1.38x 1.23x 1.02x 0.93x 0.94x 0.88x 0.99x 0.99x
EAGLE-3 1.81x 1.82x 1.62x 1.48x 1.39x 1.32x 1.38x 1.34x 1.38x
The SGLang team also tested the latency of EAGLE-3 on H100 when the target model is LLaMA-Instruct 3.1 8B and the testing dataset is MT-bench. The results are shown in Table 4.
Table 4: Latency on H100 when the target model is LLaMA-Instruct 3.1 8B and the testing dataset is MT-bench. The experiments were conducted by the SGLang team.
Method Latency (bs=1)
SGLang (w/o speculative, 1x H100) 158.34 tokens/s
SGLang + EAGLE-2 (1x H100) 244.10 tokens/s
SGLang + EAGLE-3 (1x H100) 373.25 tokens/s
4.4 EAGLE-3 in vLLM
We also conducted a study on the impact of EAGLE-3 on throughput for large batch sizes based on vLLM Kwon et al. (2023), a widely used production-grade framework, and the results on RTX3090 and LLaMA-Instruct 3.1 8B are shown in Table 5. EAGLE shows the maximum throughput improvement at a batch size of 24, while EAGLE-3 shows this at 56. This part of the experiment did not use the tree structure, the maximum chain length was set to 2, and the testing dataset was MT-Bench.
Table 5: Throughput improvement under different batch sizes on A100 and LLaMA-Instruct 3.1 8B for the MT-Bench dataset, with vLLM without speculative sampling as the baseline (1.00x).
@@ 6 Conclusion
Acknowledgement
We woud like to thank James Liu, Ke Bao, Yineng Zhang, Lianmin Zheng, Ying Sheng, and many others in the SGLang team for merging and evaluating EAGLE-3 in the SGLang environment.
@@ Appendix A Implementation Details
EAGLE-3: EAGLE-3’s draft model achieves a significantly higher acceptance rate, allowing us to increase the draft tree depth from 6 to 8 while keeping the number of nodes the same as in EAGLE-2.